find the equation of the circles which touches the line 3x+y+3=0 at (-3,6) and tangent to x+3y-7=0.
 
I AM REQUESTING ARUN TO PLEASE SOLVE THE WHOLE SUM AND NOT SIMPLY GIVE THE STEPS....

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
By family of circles
 
Eqn of circle-
(x +3)² + (y -6)² + s(3x+y+3) = 0
 
x² + y² + x(6 +3s) + y (s -12) + (45 +3s) = 0
 
Also this is a tangent to x + 3y -7 = 0
 
Hence
 
Now perpendicular length = radius thus you get value of s.